OpenStudy (anonymous):

4(2x -3) + 8(x - 4) = 2(2x + 6)

OpenStudy (anonymous):

odubo - why don't you try and answer them and them post them with your answer and we can let you know if you made an error

OpenStudy (anonymous):

8x-12+8x-32=4x+12 8x-8x-4x=12+12+32 -4x=56 x=-14

OpenStudy (anonymous):

yea i thinx mrj is ryt